# Towards the Terminator Economy: Assessing Job Exposure to AI through LLMs

### v2

@@ Abstract
AI and related technologies are reshaping jobs and tasks, either by automating or augmenting human skills in the workplace. Many researchers have been working on estimating if and to what extent jobs and tasks are exposed to the risk of being automatized by AI-related technologies. Our work tackles this issue through a data-driven approach by: (i) developing a reproducible framework that uses cutting-edge open-source large language models to assess the current capabilities of AI and robotics in performing job-related tasks; (ii) formalizing and computing a measure of AI exposure by occupation, the TEAI (Task Exposure to AI) index, and a measure of Task Replacement by AI, the TRAI index, both validated through a human user evaluation and compared with the state of the art. Our results show that the TEAI index is positively correlated with cognitive, problem-solving and management skills, while it is negatively correlated with social skills. Applying the index to the US, we obtain that about one-third of US employment is highly exposed to AI, primarily in high-skill jobs, requiring a graduate or postgraduate level of education. We also find that AI exposure is positively associated with both employment and wage growth in 2003-2023, suggesting that AI has an overall positive effect on productivity. Considering specifically the TRAI index, we find that even in high-skill occupations, AI exhibits high variability in task substitution, suggesting that AI and humans complement each other within the same occupation, while the allocation of tasks within occupations is likely to change.
@@
Keywords:
AI, Large Language models, employment, skills.
@@ 1 Introduction
The famous 1984 movie “The Terminator” is set in a dystopian future where intelligent machines, created by a military defense system known as Skynet, become self-aware and perceive humanity as a threat, initiating a war to eliminate humans. Skynet creates advanced humanoid robots called Terminators to hunt down and kill human survivors. The Terminator has advanced learning algorithms that enable it to adapt to any environment, making it a formidable antagonist for humans. The debate and concerns about the impact of AI are often framed against the backdrop of the film’s setting. This paper takes these concerns seriously by developing an AI-centered assessment of the potential exposure of different occupations to artificial intelligence. Assessing the impact of technology on the labor market is not easy, as there are several potential channels at work. As stressed by (Acemoglu and Restrepo, 2019), technology has three major effects on labor demand. The first is the productivity effect, which operates through lower production costs brought about by new technologies. The second is the displacement effect of workers by machines and the like. These two effects operate in different directions and depend on whether technology substitutes for or complements human labor; in economic jargon, they depend on the elasticity of substitution between tasks. In addition, there is a third effect of technology: the creation of new tasks and activities in which labor can be productively employed (the reinstatement effect). Indeed, if we look at history, the reinstatement effect has been a central feature of all technological revolutions, which have continually created new opportunities for labor. For the reinstatement effect to take place, technology must have an impact wider than its narrow scope, with spillover effects in sectors/areas other than those for which it was designed. In other words, technology must have the features of a general purpose technology, which, according to (Lipsey et al., 2005), are pervasiveness across the economy, ability to generate complementary innovations, and improvement over time. AI provides strong arguments for being considered a general-purpose technology due to its broad applicability, potential productivity gains, and possibility to drive further innovation. However, these characteristics of AI create a relevant measurement problem, as it is extremely difficult to identify all the channels through which it affects the economy. This paper contributes to this field by developing a methodology for assessing AI exposure using Large Language Models (LLMs), with a granular approach that analyzes exposure for each task within each occupation. Our approach allows us to improve our understanding of the effect of AI on the labor market in several areas. Methodologically, we design and implement a reproducible framework to assess to what extent existing AI and robotics technologies can perform job-related tasks, relying LLMs. Rather than assessing AI exposure through external benchmarks such as expert judgment or AI patents and innovations data, we construct an internal assessment using LLM’s own evaluation. In other words, LLMs assess their ability to perform each of the 19K job-related tasks associated with about 1k occupations as described by O*NET [^1] O*NET: Occupational Information Network, is a comprehensive database containing detailed information on hundreds of standardized and occupation-specific descriptors. It is sponsored by the U.S. Department of Labor/Employment and Training Administration. O*NET serves as a resource to provide information on skills, abilities, knowledge, work activities, and interests associated with occupations. This approach has two main advantages. On the one hand, it is very granular. Each individual task, in fact, receives a specific assessment and a motivation by each LLM. On the other hand, it is fully transparent. Not only does it allow full reproducibility, [^2] All codes are available on GitHub at https://github.com/Crisp-Unimib/Terminator-Economy but more importantly, it allows comparability over time. AI is an innovative and fast-growing field where new models are expected to significantly outperform existing ones. New LLMs can be applied to our approach to measure the extent of their improvement in task performance. Consider taxi drivers, often used as an example of a potentially exposed occupation. The current assessment of the task of driving vehicles by LLMs is of average exposure, recognizing that current technologies can replace some tasks but not others that need to be performed by humans. However, with advances in AI systems, it is likely that the same activity will be fully automated in a few years’ time. New models will therefore assign a higher score and allow the extent of the change in exposure to be measured. From an economic perspective, we develop an AI exposure measure for each O*NET occupation based on specific assessment of individual tasks . Our AI exposure index is positively correlated with cognitive, problem-solving and management skills, consistent with the evidence that AI affects management and decision-making tasks; on the other hand, AI exposure is negatively correlated with social skills, where AI has clear limitations. We also develop a version of the index which is specifically constructed for measuring task substitutability by AI and which can be used to analyze the relationship between AI exposure and AI task substitutability. This analysis is enhanced by the fact that our approach provides, for each task, a description of AI involvement which constitutes a valuable wealth of information that can be used to analyze the human-machine interaction. Bringing our measure to the data, we show that, in the US, approximately 1/3 of employment is highly exposed to AI technologies; most of this employment (88%) is in high-skill jobs, requiring graduate or postgraduate level of education. In terms of labor market outcomes, using 4-year rolling regressions, we find that AI exposure is positively associated with both employment and wage growth in the period 2003-2023, suggesting that AI has an overall positive effect on productivity. The remainder of the paper is structured as follows: section 2 discusses the related literature, section 3 describes the methodology and the construction of the AI index, section 5 presents the results; finally, section 6 concludes.
2 Background and Related Works
AI and jobs
Since the seminal paper by (Autor et al., 2003), the task approach has proved very effective in analyzing the impact of technology on jobs. It divides work activities into tasks, each of which can be performed by humans or by machines. In this way, the distinction between capital and labor tasks is more precise, flexible, and able to shift over time. Indeed, capital and machines can substitute for labor in the performance of some tasks, while complementing it in the performance of others.
@@ 2 Background and Related Works
The task approach has been applied to analyze the effect of technology and trade (offshoring) (Acemoglu and Autor, 2011), to the long run effect of technology (Consoli et al., 2023) and to skill-task interaction (Colombo et al., 2019).
@@
More recently this approach has been used to measure occupational exposure to computers and robots. In a seminal paper, (Frey and Osborne, 2017) estimated that up to 47% of jobs in the US are at risk of automation. [^3] See also (Nedelkoska and Quintini, 2018) for a similar approach. Subsequently, other attempts focused on developing measures of exposure to machine learning and robotics (Brynjolfsson and Mitchell, 2017; Acemoglu and Restrepo, 2020) and to AI (Felten et al., 2021; Webb, 2023; Eloundou et al., 2024; Pizzinelli et al., 2023).
@@
Overall, these works find an extensive share of employment exposed to AI; the specific impact on occupations varies depending on the nuances that different indicators of the impact of technology capture, i.e., whether they focus on aspects of technology that impact more routine-based activities (Frey and Osborne, 2017) or more cognitive elements (Felten et al., 2021). A common feature of these papers is the attempt to quantify AI exposure through an external benchmark, which may be expert judgment or data analysis on patents and innovations. In contrast, our approach is based on an internal assessment, where LLM systems are asked to assess the suitability of tasks for AI. This approach has two main advantages. First, it is fully transparent, with results and outcomes being fully disclosed. Second, the approach is fully reproducible. This means that whenever new generations of LLM are available, they can be used in our approach to measure the change in task exposure that they imply.
Large Language Models
LLMs are powerful computational models designed to understand and generate human-like text by harnessing vast amounts of textual data, and have taken Natural Language Processing (NLP) by storm, achieving state-of-the-art performance in many tasks (Min et al., 2023). Typically, these models are based on Transformer architecture (Vaswani et al., 2017), powered by Attention mechanism (Luong et al., 2015; Bahdanau et al., 2014) and are composed by decoder-only stack.
These models are initially trained on an Autoregressive task Radford et al. (2018), where given a sequence of words $S=(w_{1},w_{2},...,w_{n-1})$ the training objective is to maximize the log-likelihood
@@
$\sum_{i}\log P(w_{i}|w_{1},w_{2},....,w_{i-1};\theta^{T})$ where $\theta^{T}$ are the model parameters, in order to predict the next word in the sequence $\prod_{i=1}^{n}P(w_{i}\mid w_{1},...,w_{i-1})$. After being pre-trained, these models are fine-tuned for several tasks, providing examples of Natural Language Inference (Radford et al., 2018). Thanks to their capability to learn from context, known as in-context learning (Radford et al., 2019), LLMs can accomplish specific tasks with high accuracy (Zhao et al., 2023), exploiting prompt engineering methodologies such as zero-shot (Wei et al., 2021) and few-shot learning (Brown et al., 2020). In February 2025, Anthropic published a study [^4] https://www.anthropic.com/news/the-anthropic-economic-index to assess AI’s impact on jobs and tasks using O*NET, mapping millions of Claude conversations to O*NET tasks. By contrast, our approach (i) maps every O*NET occupation and task, (ii) employs a panel of three open LLMs to ensure reproducibility and minimize hallucinations, and (iii) allows the community to estimate the AI impact on other third-party taxonomies.
3 Building the AI Exposure Index
Our method can be summarized as follows: First, we obtain from O*NET the description of each task associated with each SOC occupation. Second, we apply LLMs to the task descriptions to obtain a rating of how well AI technologies can perform each task. Third, we aggregate the rating at the occupation level to obtain an AI exposure score. Finally, we apply our score to US data to assess the extent of AI exposure in the US labor market and the impact of AI on employment and wages. Figure 1 provides a graphical representation of our approach.
Figure 1: Graphical overview of the framework to compute the TEAI Index
@@ 3.1 Step 1: Compute the AI rate
In order to limit the risk of being driven by the well-known problem of LLMs called "hallucinations" (Ji et al., 2023), we design a framework involving three different LLMs aimed at identifying and limiting the false information generated, creating a consensus system between them.
Model choice
@@
To ensure the reproducibility of this work we use three of the best-performing open source models, according to performance benchmarks, available on the open LLM rankings. [^5] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard The three selected models are Mistral 7B Instruct v 0.2 (Jiang et al., 2023), openchat 3.5 0106 (Wang et al., 2023) and orca mini v3 7b (Mukherjee et al., 2023).
Prompt design
@@
The starting point is the ONET taxonomy which identifies 19281 tasks for 923 SOC occupations. [^6] We use the ONET 28.2 version released in February 2024. We formulate a five-shot prompt using the few-shot learning approach (Brown et al., 2020). We use each individual task description assigned to an occupation to ask the models how well, on a scale of 1 to 5, [^7] How well an AI system, which can be an LLM, Image Processing System or a Robot, could perform in the task on a scale of 1 to 5 where 1 stands for poor and 5 stands for excellent? the combination of different AI technologies could perform the input task and provide a discursive motivation for the evaluation. As AI technologies, we consider i) LLMs for textual data understanding, ii) Image Processing Systems for elaboration and decision-making based on visual data analysis, and iii) Robotic systems for physical execution. At the end of the prompt, the model is provided with five examples of the task to obtain more contextual and accurate results.
Model results
@@
We iterate this process for each task provided by ONET and for each model chosen, ending up with three scores and natural language motivations provided by each model. Table 1 provides an example of the results after this stage for a selection of occupations and tasks.
O*NET Task Mistral Orca-mini Open-Chat Motivation provided by Mistral Direct or coordinate the supportive services department of a business, agency, or organization. 4 4 3 Robotics can automate repetitive tasks such as data entry, document scanning, and mail sorting. Image Processing Systems can analyze documents and emails for information extraction and categorization. LLMs can handle complex tasks like scheduling, coordinating resources, and communicating with team members. Combining these AI technologies can significantly improve the efficiency and accuracy of administrative services, allowing managers to focus on more strategic tasks. Conduct classes to teach procedures to staff. 2 3 2 LLMs could potentially create instructional materials, but they lack the ability to engage and assess staff performance during classes. Image Processing Systems have no role in this task. Robotics could theoretically conduct classes through pre-recorded videos, but the human interaction and customization required for effective training make this a poor solution.
@@
Evaluating consensus among LLMs
As mentioned above, the decision to use three different models was made to limit hallucinations. In order to construct a single indicator, we took a conservative approach by assigning to each task the value of the rating with the highest frequency among the three models; if the three rates were different, we chose the lowest.
To assess the agreement between the three rates expressed by the LLMs, for each single task we compute a consensus metric (Tastle and Wierman, 2007).
Equation (1) shows the consensus calculation in which, given an input task $a_{i}$, for which three rates are expressed by three different LLMs, $LV_{k}$ denotes the rate (on a scale from 1 to 5) expressed by the $k$-th LLM, $p_{k}$ denotes the relative frequency associated with $LV_{k}$; $\mu_{LV}$ defines the weighted average of the three $LV$ rating values computed using $p_{k}$ probabilities as weights; finally $d_{LV}$ represents the range of the rating scale adopted.
Specifically, for each $k$-th rate from 1 to 3 (corresponding to the three available rates), the absolute difference between the rating value $LV_{k}$ and the weighted average $\mu_{LV}$ is determined. This difference is then normalized by dividing by the scale size $d_{LV}$. The logarithmic function $\log_{2}$ is applied in order to moderate the impact of larger discrepancies from the mean.
This process allows the three rates to be combined into a single measure of agreement. A higher consensus value indicates greater agreement between the observed rating values expressed for each task by the three different LLMs.
Analogously, to estimate the similarity between the motivations provided by the LLMs, we compute the centroid of semantic cosine similarity (Rahutomo et al., 2012), between them. The embedding vectors for the centroid computation are obtained using an open source Transformer model: as for the LLMs, we chose the Transformer model to be used in accordance with the Massive Text Embedding Benchmark (MTEB) Leaderboard by (Muennighoff et al., 2023). [^8] https://huggingface.co/spaces/mteb/leaderboard. Having English-language motivations, the choice fell on the UAE-Large-V1 https://huggingface.co/WhereIsAI/UAE-Large-V1, which represented an excellent compromise between effectiveness and efficiency, given its small size.
Table 2 shows the results following the calculation of semantic similarity and consensus between the different rates. In particular, since higher cosine similarity values reflect higher semantic similarities between the text of the LLM motivations, we expect a strong correlation between the consensus metric and cosine similarity. The heat map represented in figure 2 shows that both the values of cosine similarity and the consensus metric are extremely high, with an average close to 0.9 in both cases.
On the one hand, this suggests coherence between LLM-generated rates and the associated motivations, and on the other hand, it adds robustness to our conservative approach in choosing between different models.
The result of this process is a single score te that returns a metric from 1 to 5 measuring the extent to which AI can perform each specific task, and a quantitative indicator of the similarity between the discursive motivations generated by the models.
O*NET Occupation O*NET Task Similarity Consensus
Administrative Services Managers Direct or coordinate the supportive services department of a business, agency, or organization. 0.918 0.828
Conduct classes to teach procedures to staff. 0.935 0.828
Taxi Drivers Complete accident reports when necessary. 0.948 0.828
Vacuum and clean interiors and wash and polish exteriors of automobiles. 0.925 0.828
Table 2: The two occupation profiles and two tasks from O*NET, shown in Table 1, enriched with the centroid of semantic similarity and consensus measure.
@@ 3.2 Step 2: Compute the AI exposure
To compute occupation exposure to AI, we follow (Felten et al., 2021) and aggregate the te scores at the occupation level by weighting them by task relevance ($R$), importance ($I$), and frequency ($F$) as measured by ONET. [^9] These weights capture different aspects of the tasks. More specifically. Importance: indicates the degree to which a particular descriptor is important for the occupation. Relevance refers to the proportion of incumbents who rated the task as relevant to their job. Frequency refers to the frequency of each task within the occupation, from annual to hourly. Despite providing task descriptions, ONET does not provide rating, importance and frequency of tasks for 39 occupations. We manually assigned values for them. More specifically, for each task $j$ and occupation $i$ our AI exposure score is computed as follows:
@@
where $TE_{ij}$ identifies the metric developed in step 1 at task level, $n$ defines the number of tasks within each occupation. Each weight is scaled by its maximum to obtain equal weights. The ONET model uses different scales for Relevance (scale 1-100), Importance (scale 1-5), and Frequency (scale 1-7). We normalize the indexes to ensure equal scale across weights. Finally, the score is normalized to ensure comparability with other similar scores.

### v3

@@ Abstract
AI and related technologies are reshaping jobs and tasks, either by automating or augmenting human skills in the workplace. Many researchers have been working on estimating if and to what extent jobs and tasks are exposed to the risk of being automatized by AI-related technologies. Our work tackles this issue through a data-driven approach by: (i) developing a reproducible framework that uses cutting-edge open-source large language models to assess the current capabilities of AI and robotics in performing job-related tasks; (ii) formalising and computing a measure of AI exposure by occupation, the TEAI (Task Exposure to AI) index, and a measure of Task Replacement by AI, the TRAI index, both validated through a Human user evaluation and compared with the state-of-the-art.
Our results show that the TEAI index is positively correlated with cognitive, problem-solving, and management skills, while it is negatively correlated with social skills. Results also suggest about one-third of U.S. employment is highly exposed to AI, primarily in high-skill jobs requiring a graduate or postgraduate level of education. We also find that AI exposure is positively associated with employment and wage growth in 2003–2023, suggesting that AI has had an overall positive effect on productivity. Considering specifically the TRAI index, we find that even in high-skill occupations, AI exhibits high variability in task substitution, suggesting that AI and humans complement each other within the same occupation, while the allocation of tasks within occupations is likely to change.
@@ 1 Introduction
Artificial Intelligence (AI) provides strong arguments for being considered a general-purpose technology due to its broad applicability, potential productivity gains and possibility to drive further innovation Crafts-21; Brynjolfsson-Rock-Syverson-19; deming2025technological; WEF2025. However, these characteristics of AI create a relevant measurement problem, as it is extremely difficult to identify all the channels through which it affects the economy. Meanwhile, in late 2018, Large Language Models (LLMs) were introduced as powerful computational tools designed to understand and generate human-like text by leveraging vast amounts of data, see min2023recent. Due to their performance and adaptability as off-the-shelf models, they have become central to numerous systems capable of processing and generating text, driving applications in natural language understanding, translation, and content creation across domains.
Contribution.
This paper develops a methodology for assessing AI exposure using LLMs, with a granular approach that analyses exposure for each job task within each job occupation from two different perspectives:
(i) From a methodological point of view, we design and implement a reproducible framework to estimate to what extent existing AI and robotics technologies can perform job-related tasks relying on LLMs. Instead of assessing AI exposure through external benchmarks such as expert judgment or AI patents and innovations data, we construct an internal assessment using LLM’s own evaluation, using O*NET [^1] O*NET is the official US comprehensive taxonomy of about 1K occupations and 19K+ job-related tasks as reference occupation/task taxonomy. The result is a framework that enables individuals and organizations to develop policies for integrating AI into careers and the workforce.
(ii) From an economic perspective, we develop a measure to estimate the AI exposure and replacement index by tasks for all the occupations of the US taxonomy. We conduct an in-depth analysis of the impact of AI across occupations in the US labor market, examining job characteristics such as skill levels and educational requirements. We investigate the relationship between AI exposure and labor market outcomes like employment and wage growth over 2019–2023. This combined analysis allows us to estimate the impact of AI and technology on different occupations and economic indicators.
Reproducibility and Demo.
Codes and results are available on GitHub [^2] https://github.com/Crisp-Unimib/Terminator-Economy while an interactive online demo is available at http://terminatoreconomy.com/
2 Related Works on AI and Jobs
@@
Since the seminal paper by Autor-Levy_Murnane-2003, the task approach has been applied to analyze the effect of technology and trade (offshoring) Acemoglu-Autor-11, to the long run effect of technology Consoli-Marin-Rentocchini-Vona-2023, and to skill-task interaction Colombo-Mercorio-Mezzanzanica-2019.
@@
More recently, this approach has been used to measure occupational exposure to computers and robots. In a seminal paper Frey-Osborne-17 estimated that up to 47% of jobs in the US are at risk of automation. [^3] See also Nedelkoska-Quintini-18 for a similar approach. Subsequently, other attempts focused on developing measures of exposure to machine learning and robotics Brynjolfsson-Mitchell-17; Acemoglu-Restepo-2020 and to AI Felten-Raj-Seamans-21; Webb-23; Eloundou-Manning-Mishkin-Rock-23; Pizzinelli-Panton-Tavares-Cazzaniga-Li-23.
A common feature of those is the attempt to quantify AI exposure through an external benchmark, which may be expert judgment or data analysis on patents and innovations.
@@
In contrast, our approach is based on an internal assessment, where LLM systems are asked to assess the suitability of tasks for AI. This approach has two main advantages. First, it is fully transparent, with results and outcomes being fully disclosed. Second, the approach is fully reproducible. This means that whenever new generations of LLMs are available, they can be used in our approach to measure the change in task exposure that they imply.
@@
In February 2025, Anthropic published a study [^4] https://www.anthropic.com/news/the-anthropic-economic-index to assess AI’s impact on jobs and tasks using O*NET, mapping millions of Claude conversations to O*NET tasks. By contrast, our approach (i) maps every O*NET occupation and task, (ii) employs a panel of three open LLMs to ensure reproducibility and minimize hallucinations, and (iii) allows the community to estimate the AI impact on other third-party taxonomies.
3 Building the AI Exposure and Substitutability Indexes
Our method can be summarised as in Fig. 1.
Figure 1: Graphical overview of the framework
@@ 3.1 Step 1: Compute the AI Rate
To mitigate the risk of being driven by the well-known problem of LLMs, known as ”hallucinations” ji2023survey, we design a framework involving three different LLMs that aim to identify and limit the false information generated, thereby creating a consensus system among them.
Model choice.
@@
To ensure the reproducibility of this work, we use three of the best-performing open-source models (at the time of the experimental phase), according to performance benchmarks, available in the open LLM ranking. [^5] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard The three selected models are Mistral 7B Instruct v 0.2 jiang2023mistral, openchat 3.5 0106 wang2023openchat and orca mini v3 7b mukherjee2023orca.
Prompt design.
@@
The starting point is the ONET taxonomy which identifies 19281 tasks for 923 SOC occupations. [^6] We use the ONET 28.2 version released in February 2024. We formulate a five-shot prompt using the few-shot learning approach brown2020language. We use each individual task description assigned to an occupation to ask the models how well, on a scale of 1 to 5, [^7] How well an AI system, which can be an LLM, Image Processing System or a Robot, could perform in the task on a scale of 1 to 5 where 1 stands for poor and 5 stands for excellent? the combination of different AI technologies could perform the input task and provide a discursive motivation for the evaluation. As AI technologies, we consider i) LLMs for textual data understanding, ii) Image Processing Systems for elaboration and decision-making based on visual data analysis, and iii) Robotic systems for physical execution. At the end of the prompt, the model is provided with five examples of the task to obtain more contextual and accurate results.
Model results.
@@
We iterate this process for each task provided by ONET and for each model chosen, resulting in three scores and natural language motivations provided by each model. Table 1 provides an example of the results after this stage for a selection of occupations and tasks.
O*NET Task Mistral
Orca
mini
Open
Chat
Motivation
Mistral
Direct or
coordinate
the
supportive
services
department
of
a business
agency, or
organization.
4 4 3
Robotics can automate
repetitive tasks such as
data entry, document scanning,
and mail sorting. Image
Processing Systems can
analyze documents and emails
for information extraction and
categorization. LLMs can handle
complex tasks like scheduling,
coordinating resources, and
communicating with team
members. Combining these
AI technologies can significantly
improve the efficiency and
accuracy of administrative
services, allowing managers to
focus on more strategic tasks.
Conduct
classes to
teach
procedures
to staff.
2 3 2
LLMs could potentially
create instructional materials,
but they lack the ability to
engage and assess staff
performance during classes.
Image Processing Systems
have no role in this task.
Robotics could theoretically
conduct classes through
pre-recorded videos, but
the human interaction and
customization required for
effective training make this a
poor solution.
@@
Evaluating Consensus among LLMs.
As mentioned above, the decision to use three different models was made to limit hallucinations. In Section A.1, we describe the analysis conducted on the output of the three models and how we obtain a single indicator for each task, defined as te, a metric ranging from 1 to 5 that measures the extent to which AI can perform each specific task.
@@ 3.2 Step 2: Compute the AI Exposure
To compute occupation exposure to AI, we follow Felten-Raj-Seamans-21 and aggregate the te scores at the occupation level by weighting them by task relevance ($R$), importance ($I$), and frequency ($F$) as measured by ONET. [^8] Weights capture different aspects of the tasks. Importance: indicates the degree to which a particular descriptor is important for the occupation. Relevance refers to the proportion of incumbents who rated the task as relevant to their job. Frequency refers to the frequency of each task within the occupation, from annual to hourly. For each task $j$ and occupation $i$ our AI exposure score is computed as follows:
@@
where $TE_{ij}$ identifies the metric developed in step 1 at the task level, $n$ defines the number of tasks within each occupation. Each weight is scaled by its maximum to obtain equal weights. The ONET model uses different scales for Relevance (scale 1-100), Importance (scale 1-5), and Frequency (scale 1-7). We normalize the indices to ensure equal scale across weights. Finally, the score is normalized to ensure comparability with other similar scores.